# Dynamical stability of evolutionarily stable strategy in asymmetric games

### v1

@@ III.1 Pure Strategy
What we want to establish can be most clearly done in the simplest non-trivial setting where $m=n=2$. Generalization of the final conclusions for $m,n>2$ (see Section IV) is straightforward but involves longer algebraic steps. In any case, we would say (cf. Hofbauer and Sigmund (1998)), a strategy profile, $(\mathbf{\hat{p}},\mathbf{\hat{q}})$, is strongly stable if a mean population strategy profile, $(\sum_{i}x_{i}{\mathbf{p}}_{i},\sum_{i}y_{i}{\mathbf{q}}_{i})$, close to $(\mathbf{\hat{p}},\mathbf{\hat{q}})$, evolves under the replicator equation to converge at $(\mathbf{\hat{p}},\mathbf{\hat{q}})$. What is desired is that a strongly stable strategy be equivalent to evolutionarily stable strategy. Since the replicator equation is written using states and not strategies, we must first establish that the ESS conditions obeyed by strategies and states are equivalent.
@@
Suppose, without any loss of generality, $(\mathbf{\hat{x}},\mathbf{\hat{y}})=\left(\begin{pmatrix}1,0\end{pmatrix},\begin{pmatrix}1,0\end{pmatrix}\right)$ is an ESS as per Definition 1b. It is trivial to see that $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}\geq\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}\implies b_{11}\geq b_{21}\implies\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{1}\geq\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{1}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}\geq\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}\implies c_{11}>c_{21}\implies\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{1}\geq\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{1}$. Furthermore in case the equalities hold in the relation between terms containing states, then it is easily seen that $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y}\implies b_{12}>b_{22}$, which in turn means $\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{2}>\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{2}$ when $\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{1}=\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{1}$; similarly, we can show $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x}\implies\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{2}>\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{2}$ when $\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{1}=\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{1}$. In conclusion, if Definition 1b satisfied by the state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$, then the conditions of definition are also satisfied by the corresponding strategy $(\mathbf{p}_{1},\mathbf{q}_{1})$. It is rather repetitive to show that the conclusion remain unchanged had $(\mathbf{\hat{x}},\mathbf{\hat{y}})=\left(\begin{pmatrix}0,1\end{pmatrix},\begin{pmatrix}0,1\end{pmatrix}\right)$—corresponding strategy profile being $(\mathbf{p}_{2},\mathbf{q}_{2})$—been chosen an ESS as per Definition 1b. Furthermore, employing the these arguments, one can observe that if we had chosen to work with Definitions 2b and 3b instead of Definition 1b the state-strategy equivalence in the definition of respective ESS’s remains intact.
Now to track the evolution of the mean population strategy, the bimatrix replicator equation for the case in hand is should be best presented as
@@
We now effectively have to show that if a $(\mathbf{\hat{x}},\mathbf{\hat{y}})$, say, $\left(\begin{pmatrix}0,1\end{pmatrix},\begin{pmatrix}0,1\end{pmatrix}\right)$—equivalently, $(\mathbf{p}_{2},\mathbf{q}_{2})$—is ESS, then they are approached by neighbourhood initial condition—equivalently, $(\sum_{i}x_{i}{\mathbf{p}}_{i},\sum_{i}y_{i}{\mathbf{q}}_{i})$ reaches $(\mathbf{p}_{2},\mathbf{q}_{2})$. This, in fact, is the case can be argued as follows: Let us suppose $(x,y)=(0,0)$ denotes the ESS. We observe that near $(0,0)$, the coefficients of $(1-x)$ and $(1-y)$ terms inside the square bracket dominates and should those terms vanish, the coefficients of $x$ and $y$ terms inside the square bracket are deciding factor for the stability of $(0,0)$. Consequently, the local asymptotic stability criteria for $(x,y)=(0,0)$ explicitly comes out to be either $(a)$ $\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{2}>\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{2}$ and $\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{2}>\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{2}$ and (b) if $\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{2}=\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{2}$ and $\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{2}=\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{2}$, then $\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{1}>\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{1}$ and $\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{1}>\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{1}$. This means that ESS under Definition 1b is indeed locally asymptotically stable.
We furthermore note that had we chosen to work with Definitions 2b and 3b above instead of Definition 1b, while the state-strategy equivalence in the definition of respective ESS’s remains intact, the ESS’s no longer must be locally asymptotically stable as conditions (a) and (b) above need not be satisfied.
III.2 Mixed Strategy
@@
The state-strategy correspondence in the case of mixed states should clearly be such that evolutionarily stable state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ corresponds to evolutionarily stable strategy $(\mathbf{\hat{p}},\mathbf{\hat{q}})$ such that $\mathbf{\hat{p}}=\hat{x}\mathbf{p}_{1}+(1-\hat{x})\mathbf{p}_{2}$ and $\mathbf{\hat{q}}=\hat{y}\mathbf{q}_{1}+(1-\hat{y})\mathbf{q}_{2}$. In this case, as well, a strongly stable $(\mathbf{\hat{p}},\mathbf{\hat{q}})$ would attract $(\sum_{i}x_{i}{\mathbf{p}}_{i},\sum_{i}y_{i}{\mathbf{q}}_{i})$, close to $(\hat{\mathbf{p}},\hat{\mathbf{q}})$. For mixed state case, $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}=\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}=\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$. The former implies
@@
Substituting the matrix elements in terms of ${\sf U}_{1}$ and rearranging the resulting expression we arrive at $\mathbf{\hat{p}}\cdot{\sf U}_{1}\mathbf{\hat{q}}=\mathbf{p}\cdot{\sf U}_{1}\mathbf{\hat{q}}.$ Similarly, $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}=\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$ implies $\mathbf{\hat{q}}\cdot{\sf U}_{2}\mathbf{\hat{p}}=\mathbf{q}\cdot{\sf U}_{2}\mathbf{\hat{p}}$. In fact, similar manipulation of terms yield that $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y}$ implies $\mathbf{\hat{p}}\cdot{\sf U}_{1}\mathbf{q}>\mathbf{p}\cdot{\sf U}_{1}\mathbf{q}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x}$ implies $\mathbf{\hat{q}}\cdot{\sf U}_{2}\mathbf{p}>\mathbf{q}\cdot{\sf U}_{2}\mathbf{p}$. Summarizing, if Definition 1b satisfied by the mixed state profile $(\mathbf{\hat{x}},\mathbf{\hat{y}})$, then the conditions of definition are also satisfied by the corresponding strategies $(\mathbf{\hat{p}},\mathbf{\hat{q}})$. Similarly, such correspondence is there for Definition 2b and 3b as well.

### v2

@@ III.1 Strong Stability
What we want to establish can be most clearly done in the simplest non-trivial setting where $m=n=2$. In any case, we would say (cf. Hofbauer and Sigmund (1998)), a strategy profile, $(\mathbf{\hat{p}},\mathbf{\hat{q}})$, is strongly stable if a mean population strategy profile, $(\sum_{i}x_{i}{\mathbf{p}}_{i},\sum_{i}y_{i}{\mathbf{q}}_{i})$, close to $(\mathbf{\hat{p}},\mathbf{\hat{q}})$, evolves under the replicator equation to converge at $(\mathbf{\hat{p}},\mathbf{\hat{q}})$. What is desired is that a strongly stable strategy be equivalent to evolutionarily stable strategy. Since the replicator equation is written using states and not strategies, we must first establish that the ESS conditions obeyed by strategies and states are equivalent.
@@
Suppose, without any loss of generality, $(\mathbf{\hat{x}},\mathbf{\hat{y}})=\left(\begin{pmatrix}1,0\end{pmatrix},\begin{pmatrix}1,0\end{pmatrix}\right)$ is an ESS as per Definition 1b. It is trivial to see that $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}\geq\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}\implies b_{11}\geq b_{21}\implies\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{1}\geq\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{1}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}\geq\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}\implies c_{11}>c_{21}\implies\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{1}\geq\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{1}$. Furthermore in case the equalities hold in the relation between terms containing states, then it is easily seen that $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y}\implies b_{12}>b_{22}$, which in turn means $\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{2}>\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{2}$ when $\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{1}=\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{1}$; similarly, we can show $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x}\implies\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{2}>\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{2}$ when $\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{1}=\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{1}$. In conclusion, if Definition 1b is satisfied by the state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$, then the conditions of definition are also satisfied by the corresponding strategy $(\mathbf{p}_{1},\mathbf{q}_{1})$. It is rather repetitive to show that the conclusion remain unchanged had $(\mathbf{\hat{x}},\mathbf{\hat{y}})=\left(\begin{pmatrix}0,1\end{pmatrix},\begin{pmatrix}0,1\end{pmatrix}\right)$—corresponding strategy profile being $(\mathbf{p}_{2},\mathbf{q}_{2})$—been chosen an ESS as per Definition 1b. Furthermore, employing these arguments, one can observe that if we had chosen to work with Definitions 2b and 3b instead of Definition 1b the state-strategy equivalence in the definition of respective ESS’s remains intact.
Now, to track the evolution of the mean population strategy, the bimatrix replicator equation for the case in hand is best presented as
@@
We now effectively have to show that if an $(\mathbf{\hat{x}},\mathbf{\hat{y}})$, say, $\left(\begin{pmatrix}0,1\end{pmatrix},\begin{pmatrix}0,1\end{pmatrix}\right)$—equivalently, $(\mathbf{p}_{2},\mathbf{q}_{2})$—is ESS, then it is approached by neighbourhood initial conditions—equivalently, $(\sum_{i}x_{i}{\mathbf{p}}_{i},\sum_{i}y_{i}{\mathbf{q}}_{i})$ reaches $(\mathbf{p}_{2},\mathbf{q}_{2})$. This, in fact, is the case can be argued as follows: Here $(x,y)=(0,0)$ denotes the ESS. We observe that near $(0,0)$, the coefficients of $(1-x)$ and $(1-y)$ terms inside the square bracket dominates and should those terms vanish, the coefficients of $x$ and $y$ terms inside the square bracket are deciding factor for the stability of $(0,0)$. Consequently, comparing the dominating terms, the local asymptotic stability criteria for $(x,y)=(0,0)$ explicitly comes out to be $(a)$ $\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{2}>\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{2}$ and $\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{2}>\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{2}$, which is the case of strict NE. However, (b) if $\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{2}=\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{2}$ then $\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{1}>\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{1}$ does not ensure that all the trajectories starting in the neighbourhood of $(0,0)$ would flow towards $(0,0)$—some flows to other phase points on the line $x=0$. Hence, local asymptotic stability of $(0,0)$ is not guaranteed; the fixed point is at best Lyapunov stable, a weaker form of stability. Similar ‘weak’ stability can be seen for the remaining two cases: (c) if $\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{2}=\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{2}$, then $\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{1}>\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{1}$; and (d) if $\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{2}=\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{2}$ and $\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{2}=\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{2}$, then $\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{1}>\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{1}$ and $\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{1}>\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{1}$. This weak stability criteria is equivalent to—as one can easily see from state-strategy correspondence discussed above—condition (ii) and (iii) in Definition 1b (and-definition) which is, thus, aptly also known as weak ESS Hofbauer and Sigmund (1998).
In summary, ESS under any of the three definitions is strongly stable if it is a strict NE, since the mean population strategy converges towards $(\mathbf{p}_{2},\mathbf{q}_{2})$. However, aforementioned weak stability can only be guaranteed by the and-definition.
III.2 Pure ESS
The pure ESS can be either strict NE or non-strict NE. Let’s first discuss what happens when it is a strict NE.
III.2.1 Strict NE as ESS
One can actually find that and-, sum-, and or- definitions are equivalent in the case ESS is strict NE; and linear stability analysis shows that satisfying any of these definitions guarantees the local asymptotic stability of the corresponding fixed point. However, since this exercise will be extended to more general asymmetric games—where both inter- and intra-specific interactions are considered—in Section IV.1, we refrain from providing the proof right here to avoid repetition. The argument presented in Section IV.1 can be easily adapted to the case of bimatrix games.
Here, because it is analytically doable for bimatrix games, let us go beyond linear stability analysis for a pure ESS $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ that happens to be an isolated fixed point as well. Consider the following function:
Realizing that the two sums in $V_{1}$ represent Kullback–Leibler (KL) divergence Kullback and Leibler (1951); Cover and Thomas (2005) of $\mathbf{\hat{x}}$ from $\mathbf{x}$ and KL-divergence of $\mathbf{\hat{y}}$ from $\mathbf{y}$, evidently $V_{1}$ must be positive except at $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ where $V_{1}=0$. If we additionally can show that $\dot{V}_{1}<0$ $\forall(\mathbf{x},\mathbf{y})$ in some neighbourhood of $(\mathbf{\hat{x}},\mathbf{\hat{y}})$, then $V_{1}$ qualifies as Lyapunov function Jordan and Smith (1999); Wiggins (2003) and the ESS must be locally asymptotically stable fixed point. To this end, on taking the time derivative, we get,
$\dot{V_{1}}=-(\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}-\mathbf{x}\cdot{\sf B}\mathbf{y})-(\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf C}\mathbf{x}),$ (21)
where we have used Eq. (17). Of course, $\dot{V_{1}}<0$ when either inequalities in (5) or inequality in (10) is satisfied. Thus, any pure ESS that is an isolated fixed point, satisfying either Definition 1c or Definition 2c has $V_{1}$ as a Lyapunov function, ensuring that the corresponding fixed point is locally asymptotically stable. We mention that while a Lyapunov function for 2ESS (or-definition) is not known, the fixed point corresponding to the pure 2ESS can be shown to be locally asymptotically stable through alternate means (see Cressman (1996)).
III.2.2 Non-strict pure NE as ESS
Figure 1: Phase portraits supplementing Examples 1–6 of the main text which should be referred for further details. The red circular marker indicates the ESS (fixed point) under focus. The blue lines are trajectories under replicator equation whereas the green dashed lines are trajectories under adjusted replicator equation.
The case of pure ESS that is non-strict (or weak) NE is laden with subtleties, mainly because the corresponding fixed point happens not to be isolated. For instance, the equivalence between the two definitions (and- and sum-definitions) in the context of linear stability and Lyapunov function, discussed above, may be broken; after all, the linear stability analysis would fail as one eigenvalue should be zero and the Lyapunov function argument would not applicable since the fixed point is not isolated. Below we highlight such points through simple examples.
Example 1 (see Fig. 1a): ESS under sum-definition need not be an ESS under and-definition.
Consider the following general class of bimatrix games:
where $b_{1}>0$ and $c_{3}>0$ such that it has a non-strict pure NE, $(\hat{x},\hat{y})=(1,1)$. To confirm whether it is an ESS, we denote the probability vector corresponding to the NE as $(\mathbf{\hat{x}},\mathbf{\hat{y}})=\left(\begin{pmatrix}1,0\end{pmatrix},\begin{pmatrix}1,0\end{pmatrix}\right)$ and the neighbourhood state as $(\mathbf{x},\mathbf{y})\equiv\left(\begin{pmatrix}1-\epsilon,\epsilon\end{pmatrix},\begin{pmatrix}1-\delta,\delta\end{pmatrix}\right)$. We find that $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{{y}}-\mathbf{x}\cdot{\sf B}\mathbf{{y}}=b_{1}\epsilon+\delta\epsilon(-b_{1}+b_{2}-b_{3})$ which is positive for infinitesimally small $\epsilon$ and $\delta$, and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf C}\mathbf{x}=-\delta\epsilon c_{3}$ which is negative violating the condition of and-definition. To find whether it satisfies sum-definition of ESS, we calculate: $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{{y}}+\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{{y}}-\mathbf{y}\cdot{\sf C}\mathbf{x}=b_{1}\epsilon+\delta\epsilon(-b_{1}+b_{2}-b_{3}-c_{3})$ which being positive for infinitesimal $\epsilon$ implies $(\hat{x},\hat{y})$ satisfies sum-definition. In passing, one can check that, as has been generally established in Section II.2, it is 2ESS only under simultaneous deviation perspective.
Next, to ascertain the nature of dynamical stability of the ESS, we write the corresponding replicator equation for payoff matrices in (22):
$\dot{x}$ $=$ $x(1-x)\left[b_{2}-b_{3}+y(b_{1}-b_{2}+b_{3})\right],$ (23)
$\dot{y}$ $=$ $-y(1-y)(1-x)c_{3}.$ (24)
We can see that the fixed point $(1,1)$ can not be asymptotically stable because $\dot{y}$—being non-positive in the fixed point’s neighbourhood—stops $y$-value of any trajectory starting in the neighbourhood from approaching $1$. This demonstrates that, in the case of non-strict NE, Definition 1c and 3c (face-value perspective) are better indicator of stability than Definition 2c. This is because failing to satisfy Definition 1c or 3c, renders the fixed point to be unstable while ESS satisfying Definition 2c wrongly predicts the corresponding fixed point to be stable in the light of Theorem 1.
Example 2 (see Fig. 1b): ESS under and-definition need not be an ESS under or-definition if one adopts face-value perspective, but it is an ESS under or-definition with simultaneous deviation perspective.
Consider the following example:
${\sf B}=\begin{pmatrix}b_{1}&b_{2}\\ b_{1}&b_{3}\end{pmatrix}~{}\textrm{and}~{}{\sf C}=\begin{pmatrix}c_{1}&c_{2}\\ c_{1}&c_{3}\end{pmatrix},$ (25)
where $b_{2}>b_{3}$ and $c_{2}>c_{3}$. We can see that $(\hat{x},\hat{y})=(1,1)$ is a non-strict NE and we can again take neighbourhood state as $(\mathbf{x},\mathbf{y})\equiv\left(\begin{pmatrix}1-\epsilon,\epsilon\end{pmatrix},\begin{pmatrix}1-\delta,\delta\end{pmatrix}\right)$. We find that $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{{y}}-\mathbf{x}\cdot{\sf B}\mathbf{{y}}=(b_{2}-b_{3})\epsilon\delta$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf C}\mathbf{x}=(c_{2}-c_{3})\epsilon\delta$ which are positive for infinitesimally small $\epsilon$ and $\delta$. Therefore, the NE satisfies and-definition but it does not satisfy or-definition in face-value perspective (since for $\epsilon=0$ both the terms become zero). However, adapting simultaneous deviation perspective puts restrictions that $\epsilon\neq 0$ and $\delta\neq 0$, thereby making it an ESS under or-definition.
To analyze the dynamical stability of the ESS, we cast the corresponding replicator equations as:
$\dot{x}$ $=(b_{2}-b_{3})x(1-x)(1-y),$ (26)
$\dot{y}$ $=(c_{2}-c_{3})y(1-y)(1-x).$ (27)
It is evident that both $\dot{x}$ and $\dot{y}$ are positive, indicating that the dynamics is driven away from $x=0$ and $y=0$ by any infinitesimal perturbation. Note that since $x=1$ and $y=1$ are line of fixed points, once a trajectory reaches at either boundary, $x=1$ or $y=1$, it terminates. Furthermore, since $\frac{dy}{dx}=\frac{ky}{x}$, where $k\equiv\frac{c_{2}-c_{3}}{b_{2}-b_{3}}$, one arrives at $y=Kx^{k}$ ($K$ being an integration constant) as equation of trajectories. This result shows that only a single trajectory can terminate at $(1,1)$; others terminate at other points on the lines $x=1$ and $y=1$. This implies absence of an open neighbourhood of $(1,1)$ such that all the phase points in the neighbourhood be asymptotically attracted towards $(1,1)$. Concluding, the non-strict NE as ESS need not satisfy Theorem 1.
III.3 Mixed ESS
@@
The state-strategy correspondence—discussed in Section III.1—in the case of mixed states should clearly be such that evolutionarily stable state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ corresponds to evolutionarily stable strategy $(\mathbf{\hat{p}},\mathbf{\hat{q}})$ such that $\mathbf{\hat{p}}=\hat{x}\mathbf{p}_{1}+(1-\hat{x})\mathbf{p}_{2}$ and $\mathbf{\hat{q}}=\hat{y}\mathbf{q}_{1}+(1-\hat{y})\mathbf{q}_{2}$. In this case, as well, a strongly stable $(\mathbf{\hat{p}},\mathbf{\hat{q}})$ would attract $(\sum_{i}x_{i}{\mathbf{p}}_{i},\sum_{i}y_{i}{\mathbf{q}}_{i})$, close to $(\hat{\mathbf{p}},\hat{\mathbf{q}})$. For mixed state case, $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}=\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}=\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$. The former implies
@@
Substituting the matrix elements in terms of ${\sf U}_{1}$ and rearranging the resulting expression we arrive at $\mathbf{\hat{p}}\cdot{\sf U}_{1}\mathbf{\hat{q}}=\mathbf{p}\cdot{\sf U}_{1}\mathbf{\hat{q}}.$ Similarly, $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}=\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$ implies $\mathbf{\hat{q}}\cdot{\sf U}_{2}\mathbf{\hat{p}}=\mathbf{q}\cdot{\sf U}_{2}\mathbf{\hat{p}}$. In fact, similar manipulation of terms yield that $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y}$ implies $\mathbf{\hat{p}}\cdot{\sf U}_{1}\mathbf{q}>\mathbf{p}\cdot{\sf U}_{1}\mathbf{q}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x}$ implies $\mathbf{\hat{q}}\cdot{\sf U}_{2}\mathbf{p}>\mathbf{q}\cdot{\sf U}_{2}\mathbf{p}$. Summarizing, if Definition 1b satisfied by the mixed state profile $(\mathbf{\hat{x}},\mathbf{\hat{y}})$, then the conditions of definition are also satisfied by the corresponding strategies $(\mathbf{\hat{p}},\mathbf{\hat{q}})$. Similarly, such correspondence is there for Definitions 2b and 3b as well.